# Animate Your Thoughts: Decoupled Reconstruction of Dynamic Natural Vision from Slow Brain Activity

### v1

@@ Appendix G Broader Impacts
The pursuit to unravel and emulate the brain’s intricate visual processing systems has been a cornerstone endeavor for scholars in computational neuroscience and artificial intelligence, aiming to catalyze the evolution of brain-inspired computational models. Recent advancements in neural decoding and the reconstruction of visual stimuli from brain activity have unveiled a lot of possibilities, yet these are accompanied by profound implications that necessitate deliberate and nuanced contemplation.
A paramount concern in this domain is the privacy and confidentiality of the neural data gathered. The integrity and security of such sensitive information are of the essence, as any compromise could have severe repercussions for individuals’ personal and medical confidentiality. This underscores an indispensable requirement for robust data protection protocols to ensure the sanctity of the data and the trust of its subjects. The onus is on the research community to implement stringent safeguards, not only to avert unauthorized access but also to foster an ethical framework that respects and protects the privacy rights of all stakeholders involved.

### v2

@@ Appendix H Ethics Statement
The pursuit of unraveling and emulating the brain’s intricate visual processing systems has been a cornerstone endeavor for researchers in computational neuroscience and artificial intelligence. Recent advancements in neural decoding and the reconstruction of visual stimuli from brain activity have opened up numerous possibilities, fueling concerns about the potential harmful use cases of mind reading.
We argue that these concerns can be alleviated for two main reasons: (1) Mind reading requires brain activity recording devices with very high spatial resolution, and data acquisition systems like fMRI, which possess high spatial resolution, are not easily portable; (2) Although there are now several portable brain activity recording devices, achieving mind reading would require the subject to maintain intense focus and cooperate with the data collection process.